OpenStudy (jenniferjuice):

Choose the correct x-intercept and y-intercept of the equation x + 2y = 4. x-intercept: (0, 2), y-intercept: (4, 0) x-intercept: (2, 0), y-intercept: (4, 0) x-intercept: (0, 4), y-intercept: (2, 0) x-intercept: (4, 0), y-intercept: (0, 2)

OpenStudy (ivettef365):

substitute the x for 0 to find the y-int substitute the y for 0 to find the x-int

OpenStudy (ivettef365):

for example, to find x-int, substitute the y for 0, so you have x +2y = 4 x + 2(0) = 4 so what is x ???

OpenStudy (jenniferjuice):

um...4?

OpenStudy (jenniferjuice):

or zero?

OpenStudy (ivettef365):

right, the x-int = 4, so it is (4,0) now try to find the y-int

OpenStudy (jenniferjuice):

0,-2 ?

OpenStudy (ivettef365):

why negative??

OpenStudy (jenniferjuice):

i dont know i thought maybe

OpenStudy (ivettef365):

ok, you have x + 2y = 4 when x = 0 then 0 + 2y = 4

OpenStudy (ivettef365):

so you have 2y = 4 -- -- 2 2

OpenStudy (jenniferjuice):

y=4

OpenStudy (jenniferjuice):

y=2

OpenStudy (jenniferjuice):

2

OpenStudy (ivettef365):

ok, now you have it right, so you have x-int = (4,0) and y-int = (0,2)

OpenStudy (jenniferjuice):

so would the answer be

OpenStudy (jenniferjuice):

x-intercept: (4, 0), y-intercept: (0, 2)

OpenStudy (jenniferjuice):

?